The blood has relative density of 1.060,just slightly higher than water. If the average adult blood is 5 L, determine the weight of the blood.
1
Expert's answer
2020-09-29T09:40:30-0400

By definition, density is

"\displaystyle \rho = \frac{m}{V}"

So, "m = \rho \cdot V =1.060 \, \frac{g}{cm^3} \cdot5 \cdot 10^3 \, cm^3 = 5.3 \cdot 10^3 \,g = 5.3 \, kg"

Answer: 5.3 kg